OpenStudy (anonymous):

use the distributive property to remove the parenthesis. simplify as much as possible. 3w^3(7-6w^5) 3w^3 * 7 = 21w^3 3w^3 * (-6w^5) = -18w^8 21w^3-18w^8 Is this right?

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Which did you pick?

OpenStudy (anonymous):

We multiply: 3w^3 * 7 Which is just multiplying 7 and 3 and keeping w^3 the same. Then we multiply: 3w^3 * -6w^5 We multiply -6 and 3 together, and then we add the exponents 3 + 5. So we have: \(3 \times -6 \times w^{3 + 5}\)

OpenStudy (anonymous):

Can you do that? @Cassandra_Lea_96

OpenStudy (anonymous):

oh yeah it wasn't multiple choice or anything that's what I did I just wanted to make sure I did it right =) I broke it up and multiplied each group like you said and then I put them together and the answer I got was 21w^3-18w^8 =)

OpenStudy (anonymous):

You got it! :D

OpenStudy (anonymous):

YES! lol =D thank you and really quick the next one says to multiply and simplify and the equation is ; (y-1)(y+3) is that FOIL? =o

OpenStudy (anonymous):

Yes, it is. We multiply: y * y y * 3 -1 * y -1 * 3

OpenStudy (anonymous):

ok cool I can do it =) I just wanted to make sure it was Foil because it didn't say it on the problem =o